OpenStudy (anonymous):

Use an appropriate Half-Angle Formula to find the exact value of the expression. cos 9π 8

hartnn (hartnn):

thats cos (9pi/8)

OpenStudy (anonymous):

yea sorry

OpenStudy (anonymous):

\[(1-\cos \pi/6)/\sin \pi/6\]

OpenStudy (anonymous):

\[\cos()\frac{ 9{\pi} }{ 8 }=\cos(\frac{ 405 }{ 2 })=\cos(\frac{ 360+45 }{ 2 })=\cos(180+\frac{ 45 }{ 2 })\] \[=\cos(180)\cos(45/2)-\sin(180)\sin(45/2)\] \[=-\cos(45/2)\]

OpenStudy (anonymous):

\[\cos(45/2)=\sqrt{\frac{ \cos(45)+1 }{ 2 }}\]

OpenStudy (anonymous):

huh?